OpenStudy (anonymous):

8x2+6x-35 factor help please

OpenStudy (mathstudent55):

Want to do it again like we did? I'll guide you step by step.

OpenStudy (anonymous):

that website should help u alot

OpenStudy (mathstudent55):

First, notice there is no common factor to be pulled out.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

@britbrat4290 How exactly did your answer explain to @romanortiz65 how to answer this question?

OpenStudy (mathstudent55):

Now we multiply ac, 8 * (-35) = -280

OpenStudy (anonymous):

mathstudent55 can u help me in another problem it tells me factor in a polynomial in descending order

OpenStudy (anonymous):

i did i put the link to it ... lol

OpenStudy (mathstudent55):

Now we need two factors of -280 that add to 6. 280 = 2^3 * 5 * 7 = 20 * 14 So use 20 and -14

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

factor the trinomial x2+7x+10 write each factor as a polynomial in descending order can u help me in this one

OpenStudy (anonymous):

can you help in this problem please

OpenStudy (mathstudent55):

Now we break up the middle term using those two numbers: \(8x^2 - 14x + 20x - 35 \) Factor by parts: \( 2x(4x - 7) + 5(4x - 7) \) Now factor out the common focator of 4x - 7: \((4x - 7)(2x + 5) \)

OpenStudy (mathstudent55):

Now let's factor this triniomial: \(x^2+7x+10 \) This problem is simpler because a = 1. In other words you have only x^2, not a number times x^2. All you need to do is find 2 numbers that multiply to 10 and add to 7.

OpenStudy (anonymous):

2,10

OpenStudy (anonymous):

i mean 2,5

OpenStudy (anonymous):

so is it (x+5) (x+2)??

OpenStudy (mathstudent55):

Right. Since this is the easier case, all you do is \((x + 2)(x + 5) \)

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

can you helpp in another problem

OpenStudy (mathstudent55):

pls start a new post. i'll help

OpenStudy (anonymous):

plz close

OpenStudy (anonymous):

3X2+36x+81

OpenStudy (anonymous):

close ur questions and start a new question plz anyways im gna have to delete medals that i dont want to you keep popping up in my notifications